OpenStudy (anonymous):

if cot x = 3.2404, how would i calculate x? :s

OpenStudy (lalaly):

\[\cot(x)=\frac{1}{\tan(x)}\]

OpenStudy (lalaly):

so \[\frac{1}{tanx}=3.240\]

OpenStudy (lalaly):

no \[\tan^{-1}(\frac{1}{3.240})=x\]

OpenStudy (anonymous):

so then x would be 17.5404 right?

OpenStudy (lalaly):

yes 17.5404

OpenStudy (anonymous):

.... :S im confused now. lalaly is telling me one thing then mel123456 is telling me another... kay so if i put 1/tan(17.5404) into my calculator i get 3.2404. okay cool thanks!

OpenStudy (lalaly):

exactly:)

OpenStudy (lalaly):

me12... Was wrong